If one side of a square is increased by 4 inches and an adjacent side is multiplied by 4,the perimeter of the resulting rectangle is 3 times the perimeter of the square. find the side of the original square.
Answer by Cromlix(4381)

The original side of the square = x
One side becomes x + 4
Adjacent side = 4x
Perimeter of rectangle = 3* perimeter of square
2(x + 4) +2(4x) = 3* 4x
2x + 8 + 8x = 12x
10x - 12x = -8
-2x = -8
Multiply both sides by -1
2x = 8
x = 4 inches
Original side of square = 4 inches.
Hope this helps.
:-)